Part A Phosgene (carbonyl chloride), COCl2, is an extremely toxic gas that is used in manufacturing...

Part A

Phosgene (carbonyl chloride), COCl2, is an extremely toxic gas that is used in manufacturing certain dyes and plastics. Phosgene can be produced by reacting carbon monoxide and chlorine gas at high temperatures:

CO(g)+Cl2(g)⇌COCl2(g)

Carbon monoxide and chlorine gas are allowed to react in a sealed vessel at 459 ∘C . At equilibrium, the concentrations were measured and the following results obtained:

Gas Partial Pressure
(atm)
CO 0.740
Cl2 1.13
COCl2 0.150

What is the equilibrium constant, Kp, of this reaction?

Express your answer numerically.

Part B

The following reaction was performed in a sealed vessel at 776 ∘C :

H2(g)+I2(g)⇌2HI(g)

Initially, only H2 and I2 were present at concentrations of [H2]=4.00M and [I2]=2.30M . The equilibrium concentration of I2is 0.0500 M . What is the equilibrium constant, Kc, for the reaction at this temperature?

Express your answer numerically.
Answer #1

Solution:

Part A) For the given reaction:

CO(g) + Cl2(g) = COCl2(g)

Kp is calculated as:

Kp = Pcocl2 / Pco x Pcl2

Where, Pcocl2, Pco and Pcl2 are the partial pressures of COCl2, CO and Cl2 respectively at equilibrium.

Putting the equilibrium partial pressures, we get,

Kp = 0.150 atm / 0.740 atm x 1.13 atm

Kp = 0.179

Part B) For the given reaction, using ICE table as,

H2(g) + I2(g) = 2 HI(g)

4.00 M -----2.30 M -----0 M. (initially)

(-X) ---------(-X)---------(+2X) (change)

(4.00 -X) ---(2.30-X)------( 2X) (equilibrium)

Given that at equilibrium, [I2]= 0.050

Thus, (2.30 - X) = 0.050

X = 2.30 - 0.050 = 2.25 M

Hence, at equilibrium,

[H2] = 4.00 - 2.25 = 1.75 M

[HI] = 2 X = 2 x 2.25 = 4.50 M

[I2] = 0.050 M

Thus,

Kc = [HI]^2 / [H2] [ I2]

Kc =(4.50)^2 / (1.75) ( 0.050)

Kc = 231.43